# Scale-Dropout: Estimating Uncertainty in Deep Neural Networks Using Stochastic Scale

### v1

@@ VI-G Hardware Overhead Analysis
To assess the energy consumption of the proposed approach, we estimated the required resources for implementing a network of five layers with the Scale-Dropout method, and we assumed using 10 crossbar arrays of $256\times 256$ and 10 Spin-ScaleDrop modules to implement a LeNet-5 network. The total area needed for the implementation of the LeNet-5 topology is $0.401\,\text{mm}^{2}$ comprising the crossbar arrays and the memories. The area estimation is based on NVSim and layout measurement. Given the energy consumption of the different components of our architecture shown in Table II. We used the NVSim simulator to estimate the total energy consumption for an inference run and multiplied this value by the number of forward passes (MC run). The analysis is carried out for ten forward passes ($T=10$). The energy consumption of an inference run is shown in Table VIII compared to other FPGA and CIM implementations. For a consistent benchmark, the same metrics and dataset (MNIST) as in previous studies were used. The Scale-Dropout approach demonstrates a significant improvement in energy efficiency, reaching up to 100×100$\times$100 × higher efficiency compared to the method presented in [26]. Compared to the implementation in [27], our approach is 51×51$\times$51 × better. Furthermore, compared to the implementation based on STT-MRAM [35], the proposed approach exhibits 3.77×3.$7$ better efficiency. Finally, compared to reference [32], our approach demonstrates 4.38×4.38$\times$4.38 × greater energy efficiency.
@@
TABLE VII: Layer-wise O+overhead analysis of the proposed method in comparison to existing works SpinDrop [29] and Spatial-SpinDrop [35].
@@
Related works Technology Bit resolution Energy
@@
VII Conclusion

### v2

@@ VII-C Hardware Overhead Analysis
To assess the energy consumption of the proposed approach, we estimated the required resources for implementing a network of five layers with the Scale-Dropout method, and we assumed using $10$ crossbar arrays of $256\times 256$ and 10 Spin-ScaleDrop modules to implement a LeNet-5 network. The total area needed for the implementation of the LeNet-5 topology is $0.401\,\text{mm}^{2}$ comprising the crossbar arrays and the memories. The area estimation is based on the NVSim and layout measurement. Given the energy consumption of the different components of our architecture shown in Table II. We used the NVSim simulator to estimate the total energy consumption for an inference run and multiplied this value by the number of forward passes (MC run). The analysis is carried out for ten forward passes ($T=10$). The energy consumption of an inference run is shown in Table VIII compared to other FPGA and CIM implementations. We evaluated two topologies LeNet-5 for the MNIST dataset and VGG small for CIFAR-10. For a consistent benchmark, the same metrics as in previous studies were used. The Scale-Dropout approach significantly improves energy efficiency, reaching up to 100×100$\times$100 × higher efficiency compared to the method presented in [34]. Compared to the implementation in [37], our approach is 51×51$\times$51 × better. Furthermore, compared to the implementation based on STT-MRAM [16], the proposed approach exhibits 3.77×3.$7$ better efficiency. Finally, compared to reference [39], our approach demonstrates 4.38×4.38$\times$4.38 × greater energy efficiency. To scale up the approach, we have performed an energy consumption estimation with a VGG small topology and we report 0.29 $\mu J/Image$. Thus, energy consumption remains notably low even when considering a larger dataset such as CIFAR-10.
@@
TABLE VII: Layer-wise overhead analysis of the proposed method in comparison to existing works SpinDrop [17] and Spatial-SpinDrop [16].
@@
Related works Technology Topology Bit resolution Energy
@@
[…]
VIII Discussion
VIII-A In Distribution Uncertainty Analysis
We thoroughly analyzed the performance of the proposed method in data distribution shift and out-of-distribution data in Section VII-B, it is equally important to perform well when it receives in-distribution data. This means that correct predictions should have low uncertainty and a model should accept most of them.
In our in-distribution data analysis (Table IX), we present the accepted, rejected, TPR, TNR, and AR percentages. TPR indicates the rate of correct and accepted predictions, while TNR refers to rejected and incorrect predictions. High TPR and TNR rates are desired as they suggest that most of the accepted predictions have low uncertainty, and incorrect predictions have high uncertainty. A high AR rate also indicates that most of the correct predictions are accepted.
The VGG and ResNet-18 models, with their larger size, effectively handle the complexity of the CIFAR-10 task, showing acceptance of approximately 80% and more than 80% in both TPR and TNR, plus more than 97% in AR, confirming the efficacy of our method.
On the contrary, the smaller ResNet-20 model is not optimal for handling the complexity of CIFAR-10, leading to ’uncertainty in model architecture’ [25] and consequently to greater uncertainty in prediction. To be specific, its inference accuracy is comparatively lower $\sim 86\%$ compared to $\sim 91\%$ for the other model, since it only has 16, 32, and 64 neurons in the residual blocks. Thus, it has a lower acceptance rate (41%). That means that most predictions are uncertain and our method is also effective in quantifying ’uncertainty in the model architecture’.
Note that our classification of the predictions (OOD or ID) with our approach (see Equation LABEL:eq) is conservative and prioritizes certainty. Adjusting quantile and confidence scores (see Equation 24) can increase acceptance rates closer to inference accuracy but may decrease OOD detection rates.
TABLE IX: Analysis of the proposed method using in-distribution data, showing True Positive Rate (TPR), True Negative Rate (TNR), and Acceptance Rate (AR) for various topologies.
Topology Accept Reject TPR TNR AR
VGG 77.43% 22.57% 84.29% 81.67% 97.45%
ResNet-18 77.48% 22.52% 84.00% 81.83% 97.67%
ResNet-20 41.00% 58.98% 59.24% 95.83% 99.12%
VIII-B Corruption Robustness Analysis
The proposed method is evaluated on 15 common corruptions reported in the work (CIFAR-10-C) [64] with various topologies with and without pre-processing, as shown in Fig. X. Our approach can achieve an OOD detection rate of on average $87.06\%$, $86.10\%$ and $97.64\%$ for VGG, ResNet-18, and ResNet-20 topologies, respectively, when no pre-processing is applied.
On the other hand, when the corruption robustness dataset is pre-processed by channel-wise normalizing them, i.e., they have the same channel-wise distribution as the clean CIFAR-10 data the model expects, the corruption error drastically reduces. For example, the mean corruption error for VGG was reduced from $82.84\%$ to $49.95\%$. Consequently, the uncertainty of the predictions also reduces. Specifically, our approach achieves OOD detection rates of $58.48\%$, $56.21\%$, and $87.73\%$, respectively, for VGG, ResNet-18 and ResNet-20 topologies. Therefore, pre-processing the dataset standardizes the data and improves the corruption robustness similar to the histogram equalization results reported in the work [64].
In terms of topology, in larger networks, e.g., ResNet-18, the corruption error is relatively lower. For example, in the case of Gaussian noise, the corruption error is reduced from 86.85% in VGG to 83.52% in ResNet-18. A similar trend is observed for other datasets. However, despite the fact that the ResNet-20 model is smaller than ResNet-18, it has a relatively higher corruption error because the smaller model introduces “uncertainty in model architecture” as mentioned in the previous section.
Nevertheless, there is a direct relationship between corruption error, uncertainty, and, in turn, the OOD detection rate. In cases where the accuracy is reduced by a small margin, the model uncertainty is low, and the OOD detection rate with our approach is also low. For example, the worst-case OOD detection rate for the VGG topology is $29.53\%$. This is achieved when the accuracy is reduced by only $6.89\%$ for brightness corruption. On the other hand, the highest OOD detection rate is achieved when the accuracy is reduced by $78.88\%$ for the VGG topologies.
VIII-C Variability and scalability
[…]
TABLE X: Analysis of mean corruption errors (mCE) and mean out-of-distribution detection (mOOD) detection values of different topologies when various corruptions applied CIFAR-10 with and without pre-processing (PP). All numbers represent percentages.
Noise Blur Weather Digital
Topology PP Error mCE mOOD Gauss. Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixel JPEG
VGG No 9.55 82.84 87.06 80.57 (86.85) 80.08 (86.27) 71.6 (61.07) 91.41 (86.91) 89.99 (86.41) 88.95 (86.03) 95.24 (86.45) 76.38 (82.33) 87.52 (85.51) 97.42 (88.43) 86.23 (70.71) 84.98 (90.16) 95.26 (82.59) 87.47 (81.12) 92.85 (81.83)
Yes 49.95 58.48 68.39 (70.45) 67.96 (66.54) 78.78 (87.85) 62.75 (50.72) 59.96 (50.8) 55.78 (43.25) 58.97 (46.06) 47.0 (34.86) 56.7 (46.96) 58.09 (41.78) 29.53 (16.44) 89.84 (78.31) 48.13 (30.07) 47.38 (55.28) 47.95 (29.93)
ResNet -18 No 8.48 79.49 86.10 84.38 (83.52) 83.57 (82.03) 82.66 (84.26) 95.89 (82.13) 89.60 (83.06) 93.37 (79.46) 94.48 (81.61) 80.01 (72.81) 89.97 (80.81) 98.64 (82.34) 73.90 (61.02) 82.66 (84.26) 86.98 (78.20) 74.18 (81.28) 81.25 (75.56)
Yes 47.32 56.21 66.53 (66.71) 60.42 (65.96) 73.73 (59.13) 62.75 (50.72) 58.90 (49.82) 59.54 (44.34) 61.11 (47.05) 47.7 (32.58) 53.78 (59.25) 58.09 (41.78) 28.79 (14.61) 65.71 (74.63) 47.98 (26.66) 55.15 (53.48) 42.98 (23.02)
ResNet -20 No 13.96 74.38 97.64 99.57 (85.75) 99.44 (84.97) 84.38 (16.48) 100.0 (83.23) 99.96 (85.86) 100.0 (82.70) 100.0 (83.13) 82.66 (15.74) 99.98 (84.45) 100.0 (87.63) 99.73 (73.72) 100.0 (85.62) 100.0 (83.22) 99.05 (82.67) 99.88 (80.53)
Yes 55.61 87.73 77.74 (84.34) 81.80 (82.06) 88.09 (76.32) 96.03 (57.51) 88.46 (61.83) 94.31 (50.03) 94.50 (50.95) 85.51 (41.56) 89.11 (53.56) 93.78 (47.58) 69.94 (20.31) 99.95 (78.39) 90.30 (36.82) 83.92 (60.79) 82.46 (32.16)
VIII-D Empirical Analysis of the Posterior Distribution
We have performed an empirical evaluation of our method on the CIFAR-10 dataset and the ResNet-18 topology. We have observed that as the number of Monte Carlo samples increases, the histogram of the posterior distribution for each of the 10 classes indeed approaches a Gaussian distribution and can be considered as an approximate Gaussian distribution similar to the MC-Dropout method.
Figure 9: Per class posterior distribution of ResNet-18 topology with a Monte Carlo sample size of 1000.
IX Conclusion